# Cost-of-capital valuation with risky assets

### v1

@@ Abstract
Cost-of-capital valuation is a well-established approach to the valuation of liabilities and is one of the cornerstones of current regulatory frameworks for the insurance industry. Standard cost-of-capital considerations typically rely on the assumption that the required buffer capital is held in risk-less one-year bonds. The aim of this work is to analyze the effects of allowing investments of the buffer capital in risky assets, e.g. in a combination of stocks and bonds. In particular, we make precise how the decomposition of the buffer capital into contributions from policyholders and investors varies as the degree of riskiness of the investment increases, and highlight the role of limited liability in the case of heavy-tailed insurance risks. We present a combination of general theoretical results, explicit results for certain stochastic models and numerical results that emphasize the key findings.
@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 lays out the concrete model assumptions on which we base the considerations of the paper and establishes a number of general monotonicity results for the needed insurance premium, the invested amount of the shareholders and the overall solvency capital requirement when increasing the proportion of risky assets in the management of insurance risks. For the more particular case of normally distributed insurance and financial risks we derive explicit formulas for the involved quantities in Section 3, and for lognormally distributed quantities some semi-explicit results are obtained in Section 4. In Section 5 we then give and discuss concrete numerical illustrations of the effects of risky investments on the needed insurance premiums and solvency requirements, and Section 6 concludes. All proofs are delegated to an appendix.
@@ 2 A cost-of-capital valuation with risky assets
Note that if both $R_{0}Z_{1}$ and $X_{1}$ are nonnegative and independent, then
$\mathbb{E}[(R_{0}Z_{1})\wedge X_{1}]=\int_{0}^{\infty}\mathbb{P}(R_{0}Z_{1}>t)\mathbb{P}(X_{1}>t)dt,$ (3)
which may facilitate computations. If the investment is into a risk-less bond, then $Z_{1}\equiv 1$ which gives as a special case of the valuation formula (2)
@@
which appears as Equation (8.16) in Mildenhall and Major [21, p.199]. It also corresponds to the one-period case of [9, Eq.10] and to [1, Eq.12] in the absence of limited liability of the shareholders.
@@
As an illustration, consider a Pareto-distributed insurance risk $X_{1}$ with cumulative distribution function $F(x)=1-(x/x_{m})^{-\beta}$, $x>x_{m}>0$ and suppose that $\beta>1$, ensuring a finite mean $\mathbb{E}[X_{1}]=x_{m}\beta/(\beta-1)$. Consider a purely risk-less investment and $R_{0}$ determined by $\operatorname{VaR}_{\alpha}$,
@@
The independence assumption between $X_{1}$ and $S_{1}$ is for simplicity of exposition, but also in line with the respective independence assumption in the standard model of Solvency II for non-life business lines. The assumption on the existence of a unique solution $R_{0}$ is in fact not very restrictive. For instance, one can derive the following result (the proof of which is given in Appendix A).
@@
If $\rho$ is either $\operatorname{VaR}_{\alpha}$ or $\operatorname{ES}_{\alpha}$, $X_{1}$ and $S_{1}$ are independent and take nonnegative values only, $S_{1}$ is absolutely continuous and $\mathbb{P}(X_{1}=0)$ is sufficiently small, then there exists a unique $R_{0}\geq 0$ solving (4).
@@ 2.1 Monotonicity properties
Assume that Assumption 1 holds. For $Z_{1}$ fixed, $X_{1}\mapsto R_{0}$ is increasing with respect to increasing first-order stochastic dominance. Similarly, for $X_{1}$ fixed, $Z_{1}\mapsto R_{0}$ is decreasing with respect to increasing first-order stochastic dominance.
We need to establish some overall soundness of the functional $V_{0}$.
Proposition 2.3.
Assume that Assumption 1 holds, and $\mathbb{E}[Z_{1}]\leq 1+\eta$.
@@
    For $Z_{1}$ fixed and (1), $X_{1}\mapsto V_{0}$ is increasing with respect to increasing first-order stochastic dominance.
@@
    For $X_{1}$ fixed and (1), $Z_{1}\mapsto V_{0}$ is decreasing with respect to increasing first-order stochastic dominance.
@@
Assume that Assumption 1 holds and that $S_{1}\geq_{\text{icx}}1$. If $\widetilde{w}\geq w$, $R_{0}^{\widetilde{w}}\leq R_{0}^{w}$ and $1+\eta\geq\mu_{w}$, then $V_{0}^{\widetilde{w}}\leq V_{0}^{w}$.
@@ 3.1 Value-at-Risk
If $\gamma>\nu\Phi^{-1}(1-\alpha)$, then it may be useful to express $R_{0}$ in (8) by the equivalent expression
@@
The latter follows from the fact that
$(\mu^{2}-\sigma^{2}\Phi^{-1}(1-\alpha)^{2})(\gamma^{2}-\nu^{2}\Phi^{-1}(1-\alpha)^{2})$
is the product of the two factors
$\mu\gamma-\Phi^{-1}(1-\alpha)\sqrt{\gamma^{2}\sigma^{2}+\mu^{2}\nu^{2}-\sigma^{2}\nu^{2}\Phi^{-1}(1-\alpha)^{2}}$
and
$\mu\gamma+\Phi^{-1}(1-\alpha)\sqrt{\gamma^{2}\sigma^{2}+\mu^{2}\nu^{2}-\sigma^{2}\nu^{2}\Phi^{-1}(1-\alpha)^{2}}.$
The usefulness comes from the fact that we will consider $\gamma$ and $\nu$ fixed while allowing $\mu$ and $\sigma$ to vary as a result of considering different positions in a risky asset. $\diamond$
@@ 3.2 Expected Shortfall
4 The lognormal model
The Gaussian model setting is attractive since it provides explicit expressions, through which the effects of investment in risky assets for cost of capital valuation can directly be traced. However, the symmetry and light tails of the normal distribution do not perfectly capture features in real-world data. If we consider other models than the Gaussian, then not much can be computed explicitly. However, the lognormal model presented below is an exception: here some key quantities can be explicitly computed.
Proposition 4.1.
If $Z_{1}\sim\text{LN}(m_{z},s_{z}^{2})$ and $X_{1}\sim\text{LN}(m_{x},s_{x}^{2})$ are independent, and if $R_{0}$ is determined by $\operatorname{VaR}_{\alpha}(R_{0}Z_{1}-X_{1})=0$, then
$R_{0}=\exp\bigg(m_{x}-m_{z}+\Phi^{-1}(1-\alpha)\sqrt{s_{z}^{2}+s_{x}^{2}}\bigg)$
and
$(1+\eta)V_{0}=\exp\bigg(m_{x}-m_{z}+\Phi^{-1}(1-\alpha)\sqrt{s_{z}^{2}+s_{x}^{2}}\bigg)\big(\eta+1-\exp(m_{z}+s_{z}^{2}/2)\big)\\ \quad+\int_{0}^{\infty}\overline{\Phi}\bigg(\frac{\log t-m_{x}-\Phi^{-1}(1-\alpha)\sqrt{s_{z}^{2}+s_{x}^{2}}}{s_{z}}\bigg)\overline{\Phi}\bigg(\frac{\log t-m_{x}}{s_{x}}\bigg)dt.$ (10)
The integral in expression (10) is easily evaluated by numerical integration.
In contrast to the Gaussian model, $S_{1}$ being lognormally distributed does not imply that $Z_{1}^{w}=wS_{1}+1-w$ is lognormally distributed. Hence, Proposition 4.1 is insufficient to study $R_{0},C_{0},V_{0}$ as functions of $w$ via $Z_{1}^{w}$. If $S_{1}\sim\text{LN}(m_{s},s_{s}^{2})$ satisfies $\mathbb{E}[S_{1}]=\mu$ and $\operatorname{Var}(S_{1})=\sigma^{2}$, then $m_{s}$ and $s_{s}^{2}$ are given by
$s_{s}^{2}=\log\bigg(1+\frac{\sigma^{2}}{\mu^{2}}\bigg),\quad m_{s}=\log\bigg(\mu\bigg(1+\frac{\sigma^{2}}{\mu^{2}}\bigg)^{-1/2}\bigg).$ (11)
@@ 5.2 The Lognormal model
Let us alternatively consider the lognormal model, which is of interest for two reasons. First, it allows to see the effects of heavier right tails for the insurance risk when compared to the normal model. Secondly, the lognormal assumption on $S_{1}$ corresponds to a Black-Scholes market assumption for the risky asset, which is not uncommon in this framework. We will choose the parameters of the lognormal model in such a way that the first two moments of both $X_{1}$ and $S_{1}$ match the ones from the normal model. That is, for $\mu=1.05$, $\gamma=1$ and each choice of $\sigma$ and $\nu$ we use (11) and
@@
Figures 4 zooms into the respective values of the needed premium $V_{0}^{w}$ and also depicts the bounds that were derived in Remark 2.2. One observes that the lower bound is too coarse for practical purposes, but the upper bound is remarkably sharp in these cases. As the upper bound is obtained by ignoring the limited liability, this indicates that the value of the limited liability option is quite small (which here is due to the small value $\alpha=0.005$ and the fact that the right tail of $X_{1}-R_{0}Z_{1}$ beyond its $1-\alpha=0.995$ quantile value is not sufficiently heavy to significantly impact $V_{0}$).
Figure 5 depicts the analogue of Figure 2 under lognormality for both insurance and financial risk, in this case varying the standard deviation of the insurance risk. Again, each of these situations leads to larger overall capital requirement $R_{0}^{w}$ compared to the counterparts of Figure 2, and again the solvency capital requirement $C_{0}^{w}$ is dominating the needed insurance premium $V_{0}^{w}$ by a larger extent. Figure 6 zooms into $V_{0}^{w}$ and depicts the bounds that were derived in Remark 2.2. Again the sharpness of the upper bound is noteworthy.

### v2

@@ Abstract
Cost-of-capital valuation is a well-established approach to the valuation of liabilities and is one of the cornerstones of current regulatory frameworks for the insurance industry. Standard cost-of-capital considerations typically rely on the assumption that the required buffer capital is held in risk-less one-year bonds. The aim of this work is to analyze the effects of allowing investments of the buffer capital in risky assets, e.g. in a combination of stocks and bonds. In particular, we make precise how the decomposition of the buffer capital into contributions from policyholders and investors varies as the degree of riskiness of the investment increases, and highlight the role of limited liability in the case of heavy-tailed insurance risks. With a focus on non-life insurance, we present a combination of general theoretical results, explicit results for certain stochastic models and numerical results that emphasize the key findings.
@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 lays out the concrete model assumptions on which we base the considerations of the paper and establishes a number of general monotonicity results for the needed insurance premium, the invested amount of the shareholders and the overall solvency capital requirement when increasing the proportion of risky assets in the management of insurance risks. For the more particular case of normally distributed insurance and financial risks we derive explicit formulas for the involved quantities in Section 3. In Section 4 we then give and discuss concrete numerical illustrations of the effects of risky investments on the needed insurance premiums and solvency requirements, and Section 5 concludes. All proofs are delegated to an appendix.
@@ 2 A cost-of-capital valuation with risky assets
If the investment is into a risk-less bond, then $Z_{1}\equiv 1$ which gives as a special case of the valuation formula (2)
@@
which appears as Equation (8.16) in Mildenhall and Major [22, p.199]. It also corresponds to the one-period case of [9, Eq.10] and to [1, Eq.12] in the absence of limited liability of the shareholders.
@@
As an illustration, consider a Pareto-distributed insurance risk $X_{1}$ with cumulative distribution function
$F(x)=1-(x/x_{m})^{-\beta},\quad x>x_{m}>0$ (4)
and suppose that $\beta>1$, ensuring a finite mean $\mathbb{E}[X_{1}]=x_{m}\beta/(\beta-1)$. Consider a purely risk-less investment and $R_{0}$ determined by $\operatorname{VaR}_{\alpha}$,
@@
The independence assumption between $X_{1}$ and $S_{1}$ is for simplicity of exposition (see also Remark 3.5 and Section 5). The assumption on the existence of a unique solution $R_{0}$ is in fact not very restrictive. For instance, one can derive the following result (the proof of which is given in Appendix A).
@@
If $\rho$ is either $\operatorname{VaR}_{\alpha}$ or $\operatorname{ES}_{\alpha}$, $X_{1}$ and $S_{1}$ are independent and take nonnegative values only, $S_{1}$ is absolutely continuous and $\mathbb{P}(X_{1}=0)<1-\alpha$, then there exists a unique $R_{0}>0$ solving (3).
@@ 2.1 Monotonicity properties
We need to establish some overall soundness of the functionals $R_{0}$ and $V_{0}$.
@@
Assume that Assumption 1 holds and that $Z_{1}$ takes only nonnegative values.
@@
    For $Z_{1}$ fixed, $X_{1}\mapsto R_{0}$ is increasing with respect to increasing first-order stochastic dominance.
@@
    For $X_{1}$ fixed, $Z_{1}\mapsto R_{0}$ is decreasing with respect to increasing first-order stochastic dominance.
  • (iii)
    For $Z_{1}$ fixed and and $\mathbb{E}[Z_{1}]\leq 1+\eta$, $X_{1}\mapsto V_{0}$ is increasing with respect to increasing first-order stochastic dominance.
  • (iv)
    For $X_{1}$ fixed and $\mathbb{E}[Z_{1}]\leq 1+\eta$, $Z_{1}\mapsto V_{0}$ is decreasing with respect to increasing first-order stochastic dominance.
@@
Assume that Assumption 1 holds, that $S_{1}$ takes only nonnegative values, and that $S_{1}\geq_{\text{icx}}1$. If $\widetilde{w}\geq w$, $R_{0}^{\widetilde{w}}\leq R_{0}^{w}$ and $1+\eta\geq\mu_{w}$, then $V_{0}^{\widetilde{w}}\leq V_{0}^{w}$.
@@
The following result says that, under mild conditions, some risky investment, i.e. a small $w>0$ compared to $w=0$, is always beneficial in the sense that $R_{0}^{w}<R_{0}^{0}$ and $V_{0}^{w}<V_{0}^{0}$.
Proposition 2.5.
Let $\rho=\operatorname{VaR}_{\alpha}$ and take $S_{1}$ and $X_{1}$ to be nonnegative and independent. Assume that $\mathbb{E}[S_{1}]>1$ and that $X_{1}$ has a bounded and continuous density that is nonvanishing in a neighborhood of $R_{0}^{0}=F_{X_{1}}^{-1}(1-\alpha)$. If $w\mapsto R_{0}^{w}$ is continuously differentiable, then
$\lim_{w\to 0}\frac{d}{dw}R_{0}^{w}=\lim_{w\to 0}\frac{d}{dw}V_{0}^{w}=-R_{0}^{0}(\mathbb{E}[S_{1}]-1)<0,\quad\lim_{w\to 0}\frac{d}{dw}C_{0}^{w}=0.$
Remark 2.5.
If $X_{1}=x>0$ is degenerate, then the assumptions in Proposition 2.5 do not hold. In this case $0=\operatorname{VaR}_{\alpha}(R_{0}^{w}(wS_{1}+1-w)-x)=R_{0}^{w}(w(\operatorname{VaR}_{\alpha}(S_{1})+1)-1)+x$. If further $\operatorname{VaR}_{\alpha}(S_{1})\in(-1,0)$, then
$\lim_{w\to 0}\frac{d}{dw}R_{0}^{w}=x(\operatorname{VaR}_{\alpha}(S_{1})-1)>0.$
The conclusion is that there is no $w$ which gives $R_{0}^{w}<R_{0}^{0}$ for all possible insurance risks $X_{1}$. In order for risky investment to reduce $R_{0}^{w}$ some uncertainty about the outcome of $X_{1}$ is needed. This conclusion is consistent with Theorem 1.1 in [16], although [16] focuses on convex risk measures. $\diamond$
Remark 2.6.
We focus on cost-of-capital valuation in this paper, but we could consider replacing $C_{0}=\mathbb{E}[Z_{\text{sh}}]/(1+\eta)$ in (1) by risk-neutral valuation $C_{0}=\mathbb{E}^{\mathbb{Q}}[Z_{\text{sh}}]$ and still obtain the qualitative conclusion that some risky investment is always beneficial. Under the additional assumptions that $\mathbb{E}^{\mathbb{Q}}[S_{1}]=1$, that independence between $S_{1}$ and $X_{1}$ holds also with respect to $\mathbb{Q}$, and that $X_{1}$ has a continuous density with respect to $\mathbb{Q}$ which is nonvanishing in a neighborhood of $R_{0}^{0}$, the expression for the limit of the derivative for $V_{0}^{w}$ takes the form
$\lim_{w\to 0}\frac{d}{dw}V_{0}^{w}=R_{0}^{0}(F^{\mathbb{Q}}_{X_{1}}(R_{0}^{0})-1)(\mathbb{E}[S_{1}]-1)<0,$
where $F^{\mathbb{Q}}_{X_{1}}(R_{0}^{0})=\mathbb{Q}(X_{1}\leq R_{0}^{0})$. A proof of this fact follows along the same lines as the proof of Proposition 2.5. $\diamond$
@@ 3.1 Value-at-Risk
If $\gamma\neq\nu\Phi^{-1}(1-\alpha)$, then $R_{0}$ in (8) is equivalent to
@@
This follows by multiplying both the numerator and denominator of (8) by
$\mu\gamma-\Phi^{-1}(1-\alpha)\sqrt{\gamma^{2}\sigma^{2}+\mu^{2}\nu^{2}-\sigma^{2}\nu^{2}\Phi^{-1}(1-\alpha)^{2}},$
and factoring out $\mu^{2}-\sigma^{2}\Phi^{-1}(1-\alpha)^{2}$ from the numerator. The usefulness comes from the fact that we will consider $\gamma$ and $\nu$ fixed while allowing $\mu$ and $\sigma$ to vary as a result of considering different positions in a risky asset. $\diamond$
@@
Remark 3.5.
In this paper we restrict our attention to the setting where $S_{1}$ and $X_{1}$ are independent. If the correlation coefficient $\operatorname{Cor}(S_{1},X_{1})>0$, then risky investments would not only generate a more favorable expectation (assuming $\mu>1$) but also partly hedge the insurance liability. One may wonder whether risky investments could be advisable even in the case $\operatorname{Cor}(S_{1},X_{1})<0$. In this case the expressions for $V_{0}$ and $C_{0}$ in Proposition 3.3 remain the same, but the expression for $R_{0}$ changes. With $c:=\operatorname{Cor}(S_{1},X_{1})\leq 0$, and under the assumption that $\mu>\sigma\Phi^{-1}(1-\alpha)$ and $\gamma\neq\nu\Phi^{-1}(1-\alpha)$,
$C^{w}_{0}$ $=\frac{R^{w}_{0}\mu_{w}-\gamma}{1+\eta}(1+\delta(\alpha)),$
$V^{w}_{0}$ $=\frac{\gamma(1+\delta(\alpha))}{1+\eta}+R^{w}_{0}\frac{1+\eta-\mu_{w}(1+\delta(\alpha))}{1+\eta},$
$R^{w}_{0}$ $=\frac{\gamma^{2}-\nu^{2}\Phi^{-1}(1-\alpha)^{2}}{h(w)},$
where $\mu_{w}:=w\mu+1-w$, $\sigma_{w}:=w\sigma$, and
$h(w)$ $=\mu_{w}\gamma-c\sigma_{w}\nu\Phi^{-1}(1-\alpha)^{2}-\Big((\mu_{w}\gamma-c\sigma_{w}\nu\Phi^{-1}(1-\alpha)^{2})^{2}$
$\quad-(\mu_{w}^{2}-\sigma_{w}^{2}\Phi^{-1}(1-\alpha)^{2})(\gamma^{2}-\nu^{2}\Phi^{-1}(1-\alpha)^{2})\Big)^{1/2}.$
The only difference from the case $c=0$ is that a term involving $c$ appears twice in the expression for $h(w)$. Straightforward computations give
$\frac{dR^{w}_{0}}{dw}(0)$ $=-(\gamma^{2}-\nu^{2}\Phi^{-1}(1-\alpha))\frac{h^{\prime}(0)}{h^{2}(0)}$
$=-(\gamma+\nu\Phi^{-1}(1-\alpha))(\mu-1+c\sigma\Phi^{-1}(1-\alpha)).$
We conclude that as long as $\mu>1-c\sigma\Phi^{-1}(1-\alpha)$, small positions $w>0$ in the risky asset will still imply $R_{0}^{w}<R_{0}^{0}$. $\diamond$
@@ 4.1 The Gaussian model
Finally, at first sight it may seem surprising that the increase of $R_{0}^{w}$ for increasing standard deviation $\nu$ of the insurance risk is almost exclusively swallowed by the increase of $C_{0}^{w}$ and leaves the premium $V_{0}^{w}$ virtually unchanged. The explanation is that the increased standard deviation (for fixed $\mathbb{E}[X_{1}]$) raises the value of the limited liability option. Concretely, for any absolutely continuous random variable $X_{1}$ (not only Gaussian!), it follows from (2) that
$C_{0}^{w}=R_{0}^{w}\frac{\mathbb{E}[Z_{1}]}{1+\eta}-\frac{1}{1+\eta}\mathbb{E}[(R_{0}^{w}Z_{1}\wedge X_{1})]$
Therefore
$\frac{\partial C_{0}^{w}}{\partial R_{0}^{w}}$ $=\frac{\mathbb{E}[Z_{1}]}{1+\eta}-\frac{1}{1+\eta}\mathbb{E}[Z_{1}\mathds{1}_{\{R_{0}^{w}Z_{1}<X_{1}\}}]$
$=\frac{\mathbb{E}[Z_{1}]}{1+\eta}-\frac{1}{1+\eta}\mathbb{E}[Z_{1}\mid R_{0}^{w}Z_{1}<X_{1}]\mathbb{P}(R_{0}^{w}Z_{1}<X_{1})$
With $\rho=\operatorname{VaR}_{\alpha}$ we then get
$\frac{\partial C_{0}^{w}}{\partial R_{0}^{w}}$ $=\frac{\mathbb{E}[Z_{1}]}{1+\eta}-\frac{1}{1+\eta}\mathbb{E}[Z_{1}\mid R_{0}^{w}Z_{1}<X_{1}]\alpha$
$\geq\frac{\mathbb{E}[Z_{1}]}{1+\eta}-\frac{\mathbb{E}[Z_{1}]\alpha}{1+\eta}=\frac{\mathbb{E}[Z_{1}](1-\alpha)}{1+\eta}.$ (10)
As $\mathbb{E}\left[Z_{1}\right]$ is close to $1+\eta$ and $\alpha$ is very small, most of the increase of the required capital is indeed absorbed by the shareholders’ contribution $C_{0}^{w}$ (see also Remark 2.3).
@@ 4.2 The Lognormal model
Let us alternatively consider the lognormal model, which is of interest for two reasons. First, it allows to see the effects of heavier right tails for the insurance risk when compared to the normal model. Secondly, the lognormal assumption on $S_{1}$ corresponds to a Black-Scholes market assumption for the risky asset, which is not uncommon in this framework. We will choose the parameters of the lognormal model in such a way that the first two moments of both $X_{1}$ and $S_{1}$ match the ones from the normal model. That is, for $\mu=1.05$, $\gamma=1$ and each choice of $\sigma$ and $\nu$ we use
$s_{s}^{2}=\log\bigg(1+\frac{\sigma^{2}}{\mu^{2}}\bigg),\quad m_{s}=\log\bigg(\mu\bigg(1+\frac{\sigma^{2}}{\mu^{2}}\bigg)^{-1/2}\bigg)$
and
@@
Figures 4 zooms into the respective values of the needed premium $V_{0}^{w}$ and also depicts the bounds that were derived in Remark 2.2 (the upper bound is very tight, as a further magnification in the plots indicates). One observes that the lower bound is too coarse for practical purposes, but the upper bound is remarkably sharp in these cases. As the upper bound is obtained by ignoring the limited liability, this indicates that the value of the limited liability option is quite small (which here is due to the small value $\alpha=0.005$ and the fact that the right tail of $X_{1}-R_{0}Z_{1}$ beyond its $1-\alpha=0.995$ quantile value is not sufficiently heavy to significantly impact $V_{0}$).
Figure 5 depicts the analogue of Figure 2 under lognormality for both insurance and financial risk, in this case varying the standard deviation of the insurance risk. Again, each of these situations leads to larger overall capital requirement $R_{0}^{w}$ compared to the counterparts of Figure 2, and again the solvency capital requirement $C_{0}^{w}$ absorbs most of that increase, whereas the premium $V_{0}^{w}$ stays almost unchanged for increasing standard deviation. The latter is again due to the increased value of the limited liability option, cf. (10) and also Example 2.1 for the case $w=0$.
Figure 6 zooms into $V_{0}^{w}$ and depicts the bounds that were derived in Remark 2.2. Again the sharpness of the upper bound is noteworthy.